# Cross-Modal Learning for Anomaly Detection in Fused Magnesium Smelting Process: Methodology and Benchmark

### v1

@@ VI Conclusion and Limitations
In this paper, we introduce a novel Transformer, dubbed FmFormer, through the lens of cross-modal learning to enhance anomaly detection in a scenario of fused magnesium smelting process. In the proposed FmFormer, a multiscale tokenization module is developed to handle the problem of large dimensionality gap between the two modalities of video and three-phase alternating current. On this basis, a cross-modal Transformer encoder is employed to alternatively explore the internal features of each modality and the correlation features across modalities. Through a multi-head decoder, our FmFormer is able to preform class-level anomaly prediction and pixel-level anomaly region detection. Interestingly, despite the poor detection accuracy when using a single current modality, the comprehensive performance of the model can still be improved through a simple linear fusion. Furthermore, by taking advantage of the cross-modal learning, the proposed method achieves an accurate anomaly detection under extreme interferences such as current fluctuation and visual occlusion from heavy water mist. To demonstrate the effectiveness of the FmFormer, we present the first cross-modal benchmark for anomaly detection of fused magnesium smelting processes, which possesses synchronously acquired video-current data and pixel-level labels. We hope it will be helpful to the research community of cross-modal learning in industrial scenarios.
Limitations. Promising directions not discussed in this work include joint training using a mixture of image-current cross-modality, video-current cross-modality, and unimodal inputs, which may involve adaptive sampling (tokenization) depending on the input modes. Adaptation to different input modes may help to solve the problem of temporary failure of a certain modality in practical applications. Another limitation is that only current-to-video tokens stemmed from the dilated tokenization and standard tokenization are used for the dense prediction. Despite the explicit interaction of the two modalities in the Transformer encoder, the performance of dense prediction could still be improved if video-to-current tokens and class tokens are used.

### v2

@@ VI Conclusion and Discussion
In this paper, we introduce a novel Transformer, dubbed FmFormer, through the lens of cross-modal learning to enhance anomaly detection in fused magnesium smelting process. In the proposed FmFormer, a multiscale tokenization module is developed to handle the problem of large dimensionality gap between modalities. A cross-modal Transformer encoder is then employed to alternatively explore the internal features of each modality and the correlation features across modalities. Through a multi-head decoder, our FmFormer is able to preform class-level anomaly prediction and pixel-level anomaly region detection. Interestingly, despite the poor detection accuracy when using a single current modality, the comprehensive performance of the model can still be improved through a simple linear fusion. Furthermore, by taking advantage of the cross-modal learning, the proposed method achieves an accurate anomaly detection under extreme interferences such as current fluctuation and visual occlusion from heavy water mist. To demonstrate the effectiveness of the FmFormer, we present the first cross-modal benchmark for anomaly detection of fused magnesium smelting processes, which possesses synchronously acquired video-current data and pixel-level labels. We hope it will be helpful to the research community of cross-modal learning in complex industrial scenarios.
Limitations and future works. Our FmFormer is optimized for temporally aligned inputs (i.e., synchronized data), but achieving strict alignment among multiple modalities in real-world industrial applications can be challenging. This may be addressed by leveraging the global perception capabilities of attention mechanisms along the temporal dimension, paired with a learning paradigm for unaligned modalities. Additionally, network communication instability may lead to missing data points or even inaccessible modalities. Implementing a masked cross-modal modeling approach with adaptive modality switching could help, along with joint training that combines image-current, video-current, and unimodal inputs. Currently, our implementation relies solely on current-to-video tokens from dilated and standard tokenization for dense prediction. To further enhance performance, one could also incorporate video-to-current tokens and class tokens.